# A Nonabelian Hodge Correspondence for Principal Bundles in Positive Characteristic

### v1

@@ 4.2. G-bundles on Root Stacks
We consider a special class of tame smooth Deligne–Mumford stacks given in [MO05, Cad07], which are called Cadman–Vistoli root stacks in [Sim11]. We are mostly interested in these root stacks in the rest of this paper. Let $X$ be a smooth variety with a reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$. We equip each irreducible component $D_{i}$ with a positive integer $d_{i}$ coprime to $p$. Denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of these integers.
@@
Given the data $(X,D,\boldsymbol{d})$ above, there exists a tame smooth Deligne–Mumford stack $\pi:\mathscr{X}\rightarrow X$ together with a reduced normal crossing divisor $\widetilde{D}=\sum\limits_{i=1}^{s}\widetilde{D}_{i}\subseteq\mathscr{X}$ such that
@@
We use the notation $\mathscr{X}_{(X,D,\boldsymbol{d})}$ for this stack and call it a root stack in this paper.
@@
In the following, if there is no ambiguity, we use the notation $\mathscr{X}$ for the root stack $\mathscr{X}_{(X,D,\boldsymbol{d})}$ given in Theorem 4.3. With the same notation as above, let $\eta_{i}$ be the generic point of $D_{i}$. We define a discrete valuation ring $R_{i}:=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and $\widehat{K}_{i}$ the corresponding completions. More generally, if $R$ is a local ring, we always use the notation $\widehat{R}$ for the completion and $\widehat{K}$ for the fraction field. There exists a discrete valuation ring $\widetilde{R}_{i}$ such that
@@
Furthermore, let $w_{i}$ (resp. $z_{i}$) be a generator of the maximal ideal of $\widetilde{R}_{i}$ (resp. $R_{i}$), and then the isomorphism $\mathscr{X}\times_{X}{\rm Spec}\,R_{i}\cong[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ is induced by a morphism $R_{i}\rightarrow\widetilde{R}_{i}$ such that $z_{i}$ is sent to $w_{i}^{d_{i}}$. Based on this property, we give the definition of type of $G$-bundles $\mathscr{E}$ on $\mathscr{X}$. The scheme $\mathscr{E}_{i}:=\mathscr{E}\times_{\mathscr{X}}[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ can be regarded as a $\mu_{d_{i}}$-equivariant $G$-bundle on ${\rm Spec}\,\widetilde{R}_{i}$. Since $\widetilde{R}_{i}$ is a discrete valuation ring, a $\mu_{d_{i}}$-action on $\mathscr{E}_{i}$ is given by a representation $\rho_{i}:\mu_{d_{i}}\rightarrow T$ (see [BS15, §2.2.4] or [BP22, §5] for instance).
@@
In the following, we give an equivalent description of $\mathscr{E}$ via patching local data and the main technique we use comes from [BLR90, chapter 6]. It is natural to see that a $G$-bundle $\mathscr{E}$ gives a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
@@
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
is an isomorphism for each $i$. Equivalently, the isomorphism $\Theta_{i}$ can be regarded as an element in $G(\widehat{\widetilde{K}}_{i})$ such that
$\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i}).$
@@
$G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}:=\{\Theta_{i}\in G(\widehat{\widetilde{K}}_{i})\,|\,\Theta_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\Theta_{i}(w_{i})\}.$
@@
$G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}:=\{\Psi_{i}\in G(\widehat{\widetilde{R}}_{i})\,|\,\rho_{i}(\gamma_{i})\Psi_{i}(w_{i})\rho_{i}(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i})\}.$
We consider the other direction that given a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$, where
  • $\mathscr{E}_{0}$ is a $G$-bundle on $\mathscr{X}\backslash\widetilde{D}$,
  • ^$\widehat{\mathscr{E}}_{i}$ is a $G$-bundle on $[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]$,
  • $\Theta_{i}:\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\rightarrow\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$ is an isomorphism,
we want to construct a unique $G$-bundle (up to isomorphism) on $\mathscr{X}$. If $D_{i}$ are pairwise disjoint, then we can obtain a $G$-bundle on $\mathscr{X}$ by applying the argument of [BLR90, 6.2 Proposition D.4] for root stacks. In the general case, the data $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\Theta_{i})$ is not enough to give a $G$-bundle on $\mathscr{X}$ and the main problem comes from the intersections $D_{i}\cap D_{j}$. In order to obtain a $G$-bundle on $\mathscr{X}$, the isomorphisms $\{\Theta_{i},1\leq i\leq s\}$ should be “compatible” on the intersections. In the following, we give a precise description of the compatibility condition.
If the intersection of $D_{i}$ and $D_{j}$ is not empty, we define an algebraic cycle $D_{ij}=D_{i}\cdot D_{j}$ and denote by $\eta_{ij}$ its generic point. Let $R_{ij}:=\mathcal{O}_{X,\eta_{ij}}$ be the localization with fraction field $K_{ij}$. Similarly, there exists a ring $\widetilde{R}_{ij}$ such that $\mathscr{X}\times_{X}{\rm Spec}\,R_{ij}\cong[{\rm Spec}\,\widetilde{R}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]$. For each pair $i,j$, if there is an isomorphism
$\Theta_{ij}:\widehat{\mathscr{E}_{i}}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]\xrightarrow{\cong}\widehat{\mathscr{E}_{j}}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]$
such that $(p_{0i}\circ\Theta_{i})(s_{i})=(p_{0j}\circ\Theta_{j})(s_{j})$ if and only if $(\Theta_{ij}\circ p_{ij})(s_{i})=p_{ji}(s_{j})$ for any section $s_{i}$ (resp. $s_{j}$) of ^$\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$ (resp. ^$\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{j}/\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]$), then we say that the tuple (,^$(\mathscr{E},\widehat{\mathscr{E}}_{i},\Theta_{i})$ is compatible.
${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]}$${\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]}$^${\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{j}/\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{j}/\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$^${\widehat{\mathscr{E}}_{j}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{ij}/\mu_{d_{i}}\times\mu_{d_{j}}]}$$\scriptstyle{p_{0i}}$$\scriptstyle{p_{0j}}$$\scriptstyle{\Theta_{i}}$$\scriptstyle{p_{ij}}$$\scriptstyle{\Theta_{j}}$$\scriptstyle{p_{ji}}$$\scriptstyle{\Theta_{ij}}$
It is routine to check that this will give a $G$-bundle on $\mathscr{X}$. The above discussion gives the following lemma.
Lemma 4.6.
A $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ is equivalent to a compatible tuple (,~$(\mathscr{E},\widetilde{\mathscr{E}}_{i},\Theta_{i})$.
Remark 4.7.
In the case of curves, $\widetilde{D}$ is a set of distinct points. Thus, the intersection $\widetilde{D}_{i}\cap\widetilde{D}_{j}$ is always empty. In this case, isomorphisms
$\Theta_{i}:\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
give all information of a $G$-bundle $\mathscr{E}$. Moreover, isomorphism classes of $G$-bundles on $\mathscr{X}$ of type $\boldsymbol{\rho}$ are in one-to-one correspondence with the double coset
@@
and this is an adelic description of $G$-bundles in the case of curves [BS15, §3].
@@ 4.3. Logarithmic G-Higgs Bundles and Logarithmic Flat G-bundles
In the previous subsection, a $G$-bundle on $\mathscr{X}\cong[Y/R]$ is regarded as one on $Y$ equipped with an isomorphism $\tau$. In this subsection, we define logarithmic $G$-Higgs bundles and logarithmic flat $G$-bundles on $\mathscr{X}=\mathscr{X}_{(X,D,\boldsymbol{d})}$ in the same way as what we defined on varieties in §2.1.
@@
A logarithmic $G$-Higgs bundle on $\mathscr{X}$ is a pair $(\mathscr{E},\theta)$, where $\mathscr{E}$ is a $G$-bundle on $\mathscr{X}$ and $\theta:\mathscr{X}\rightarrow\mathscr{E}(\mathfrak{g})\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is integrable. A logarithmic flat $G$-bundle on $\mathscr{X}$ is a pair $(\mathscr{V},\nabla)$, where $\mathscr{V}$ is a $G$-bundle $\mathscr{X}$ and $\nabla:\mathcal{O}_{\mathscr{V}}\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is an integrable connection.
@@
Given $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$, we obtain a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ via a logarithmic version of the inverse Cartier transform given in Theorem 4.2. For convenience, we regard $(\mathscr{E},\theta)$ (resp. $(\mathscr{V},\nabla)$) as equivariant logarithmic $G$-Higgs bundles $(E,\theta)$ (resp. logarithmic flat $G$-bundles $(V,\nabla)$) on ${\rm Spec}(k[y_{1},\dots,y_{d}])$. Thus, following the construction of (3.2) in the proof of Theorem 3.1, we have
@@ 4.4. Stability Condition
In this subsection, we make a further assumption on the root stack $\mathscr{X}$ that it is projective. In other words, the coarse moduli space $X$ is projective. We fix an ample line bundle $H$ on $X$ and a locally free sheaf $\mathscr{L}$ on $\mathscr{X}$. The modified Hilbert polynomial $P_{\mathscr{L}}(\mathscr{F},m)$ is defined as follows:
$P_{\mathscr{L}}(\mathscr{F},m)=\chi(\mathscr{X},\mathscr{F}\otimes\mathscr{L}\otimes\pi^{*}(H^{m})),\,m\gg 0.$
When $\mathscr{F}$ is pure of dimension $d$, the function $P_{\mathscr{L}}(\mathscr{F},m)$ can be written as
$P_{\mathscr{L}}(\mathscr{F},m)=\sum_{i=0}^{d}\alpha_{\mathscr{L},i}(\mathscr{F})\frac{m^{i}}{i!},$
and the $\mathscr{L}$-degree of $\mathscr{F}$ is given as
$\frac{\deg_{\mathscr{L}}\mathscr{F}}{r}=\frac{\alpha_{\mathscr{L},d-1}(\mathscr{F})}{re}-\frac{\alpha_{\mathscr{L},d-1}(\mathcal{O}_{\mathscr{X}})}{e},$
where $r$ is the rank of $\mathscr{F}$ and $e$ is the rank of $\mathscr{L}$ [Nir09, Proposition 3.18]. As pointed out in [Nir09, Corollary 7.10], with a good choice of $\mathscr{L}$, the degree $\deg_{\mathscr{L}}\mathscr{F}$ is equal to the parabolic degree of the corresponding parabolic sheaf on $X$. We also refer the reader to [Bor07] for more details.
Remark 4.14.
If $\mathscr{X}$ is a global quotient, the locally free sheaf $\mathscr{L}$ is a generating sheaf [Kre09, Theorem 5.3]. Moreover, generating sheaves are very helpful to construct the moduli spaces of coherent sheaves and Higgs sheaves on Deligne–Mumford stacks (see [OS03, Nir09, Sun20, Sun22] for instance).
We define stability conditions for logarithmic $G$-Higgs bundles and logarithmic flat $G$-bundles on $\mathscr{X}$ similar as those in §2.3.
@@
A logarithmic $G$-Higgs bundle $(\mathscr{E},\theta)$ is $R$-semistable (resp. $R$-stable), if
  • for any proper parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\theta$-compatible reduction of structure group $\sigma:\mathscr{X}\rightarrow\mathscr{E}/P$,
@@
Denote by ${\rm HIG}^{s(s)}_{p-1}(\mathscr{X}_{\rm log}/k,G)\subseteq{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$ the subcategory of $R$-(semi)stable nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$.
A logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ is $R$-semistable (resp. $R$-stable), if
  • for any pparabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ acting trivially on the center;
  • for any $\nabla$-compatible reduction of structure group $\varsigma:\mathscr{X}\rightarrow\mathscr{V}/P$,
we have
[…]
Let ${\rm MIC}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ be the subcategory of ${\rm MIC}_{p-1}(\mathscr{X}_{\rm log}/k,G)$ consisting of $R$-(semi)stable nilpotent logarithmic flat $G$-bundles of exponent $\leq p-1$ on $\mathscr{X}$.
Similar to the setup in §3.3, we define $\mathscr{L}^{\prime}:=\pi^{*}_{\mathscr{X}/k}\mathscr{L}$ and $H^{\prime}:=\pi^{*}_{X/k}H$. Then the degree of a pure sheaf $\mathcal{F}^{\prime}$ on $\mathscr{X}^{\prime}$ is given with respect to $H^{\prime}$ and $\mathscr{F}^{\prime}$. It is easy to check that $\deg_{\mathscr{L}^{\prime}}\mathscr{F}^{\prime}=p\deg_{\mathscr{L}}\mathscr{F}$. Then, the stability condition of Higgs bundles on $\mathscr{X}^{\prime}$ is defined by $H^{\prime}$ and $\mathscr{F}^{\prime}$, while the stability condition of flat bundles on $\mathscr{X}$ is given with respect to $H$ and $\mathscr{F}$.

### v2

@@ 4.2. G-bundles on Root Stacks
We consider a special class of tame smooth Deligne–Mumford stacks studied in [Bor07, MO05, Cad07, Sim11], which are called root stacks in this paper. We first review an existence theorem about such root stacks.
@@
Let $X$ be a smooth variety with a reduced normal crossing divisor $D=\sum\limits_{i=1}^{s}D_{i}$. We equip each irreducible component $D_{i}$ with a positive integer $d_{i}$ coprime to $p$. Denote by $\boldsymbol{d}=\{d_{i},1\leq i\leq s\}$ the collection of these integers. Then there exists a tame smooth Deligne–Mumford stack $\pi:\mathscr{X}\rightarrow X$ together with a reduced normal crossing divisor $\widetilde{D}=\sum\limits_{i=1}^{s}\widetilde{D}_{i}\subseteq\mathscr{X}$ such that
@@
This stack is denoted by $\mathscr{X}_{(X,D,\boldsymbol{d})}$ and is called a root stack.
Remark 4.4.
By Kawamata covering lemma [Kaw81, Theorem 17], the data $(X,D,\boldsymbol{d})$ determines a Galois covering $\pi:Y\rightarrow X$ with Galois group $\Gamma$, and thus $\mathscr{X}_{(X,D,\boldsymbol{d})}\cong[Y/\Gamma]$. Note that although we work on positive characteristic, the lemma is still true under the tameness condition (see [BP22, §12] for instance). Under this isomorphism, we regard $G$-bundles on $\mathscr{X}_{(X,D,\boldsymbol{d})}$ as $\Gamma$-equivariant $G$-bundles on $Y$, and we use the terminology $(\Gamma,G)$-bundles on $Y$ for convenience.
@@
In the following, if there is no ambiguity, we use the notation $\mathscr{X}$ for the root stack $\mathscr{X}_{(X,D,\boldsymbol{d})}$ given in Theorem 4.3. With the same notation as above, let $\eta_{i}$ be the generic point of $D_{i}$. We define a discrete valuation ring $R_{i}:=\mathcal{O}_{X,\eta_{i}}$ with fraction field $K_{i}$. Denote by $\widehat{R}_{i}$ and $\widehat{K}_{i}$ the corresponding completions. More generally, if $R$ is a local ring, we always use the notation $\widehat{R}$ for the completion with $\widehat{K}$ the fraction field. There exists a discrete valuation ring $\widetilde{R}_{i}$ such that
@@
Furthermore, let $w_{i}$ (resp. $z_{i}$) be a generator of the maximal ideal of $\widetilde{R}_{i}$ (resp. $R_{i}$), and then the isomorphism $\mathscr{X}\times_{X}{\rm Spec}\,R_{i}\cong[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ is induced by a morphism $R_{i}\rightarrow\widetilde{R}_{i}$ such that $z_{i}$ is sent to $w_{i}^{d_{i}}$. Based on this property, we give the definition of type of $G$-bundles $\mathscr{E}$ on $\mathscr{X}$. The scheme $\mathscr{E}_{i}:=\mathscr{E}\times_{\mathscr{X}}[{\rm Spec}\,\widetilde{R}_{i}/\mu_{d_{i}}]$ can be regarded as a $(\mu_{d_{i}},G)$-bundle on ${\rm Spec}\,\widetilde{R}_{i}$. Since $\widetilde{R}_{i}$ is a discrete valuation ring, a $\mu_{d_{i}}$-action on $\mathscr{E}_{i}$ is given by a representation $\rho_{i}:\mu_{d_{i}}\rightarrow T$ (see [BS15, §2.2.4] or [BP22, §5] for instance).
@@
Remark 4.7.
Now we consider the local structure. A $G$-bundle $\mathscr{E}$ on $\mathscr{X}$ gives a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\widehat{\Theta}_{i})$, where
@@
$\widehat{\Theta}_{i}:\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
is an isomorphism for each $i$. We want to remind the reader that such a tuple $(\mathscr{E}_{0},\widehat{\mathscr{E}}_{i},\widehat{\Theta}_{i})$ cannot determine a $G$-bundle on $\mathscr{X}$ in the higher dimensional case.
The isomorphism $\widehat{\Theta}_{i}$ can be equivalently regarded as an element in $G(\widehat{\widetilde{K}}_{i})$ such that
$\widehat{\Theta}_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\widehat{\Theta}_{i}(w_{i}).$
@@
$G(\widehat{\widetilde{K}}_{i})^{\rho_{i}}:=\{\widehat{\Theta}_{i}\in G(\widehat{\widetilde{K}}_{i})\,|\,\widehat{\Theta}_{i}(\gamma_{i}\cdot w_{i})=\rho_{i}(\gamma_{i})\widehat{\Theta}_{i}(w_{i})\}.$
@@
$G(\widehat{\widetilde{R}}_{i})^{\rho_{i}}:=\{\widehat{\Psi}_{i}\in G(\widehat{\widetilde{R}}_{i})\,|\,\rho_{i}(\gamma_{i})\widehat{\Psi}_{i}(w_{i})\rho_{i}(\gamma_{i})^{-1}=\Psi_{i}(\gamma_{i}\cdot w_{i})\}.$
In the case of stacky curves, $\widetilde{D}$ is a set of distinct points. Isomorphisms
$\widehat{\Theta}_{i}:\mathscr{E}_{0}\times_{\mathscr{X}\backslash\widetilde{D}}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]\xrightarrow{\cong}\widehat{\mathscr{E}}_{i}\times_{[{\rm Spec}\,\widehat{\widetilde{R}}_{i}/\mu_{d_{i}}]}[{\rm Spec}\,\widehat{\widetilde{K}}_{i}/\mu_{d_{i}}]$
give all information of a $G$-bundle $\mathscr{E}$, and isomorphism classes of $G$-bundles of type $\boldsymbol{\rho}$ on $\mathscr{X}$ are in one-to-one correspondence with the double coset
@@
and this is an adelic description of $G$-bundles in the case of curves [BS15, §3]. This result holds in higher dimensional case when the normal crossing divisor $D$ is smooth.
@@ 4.3. Logarithmic G-Higgs Bundles and Logarithmic Flat G-bundles
A logarithmic $G$-Higgs bundle on $\mathscr{X}$ is a pair $(\mathscr{E},\theta)$, where $\mathscr{E}$ is a $G$-bundle on $\mathscr{X}$ and $\theta:\mathscr{X}\rightarrow\mathscr{E}(\mathfrak{g})\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is integrable. A logarithmic flat $G$-bundle on $\mathscr{X}$ is a pair $(\mathscr{V},\nabla)$, where $\mathscr{V}$ is a $G$-bundle $\mathscr{X}$ and $\nabla:\mathcal{O}_{\mathscr{V}}\rightarrow\mathcal{O}_{\mathscr{V}}\otimes\Omega_{\mathscr{X}}({\rm log}\,\widetilde{D})$ is an integrable $G$-connection.
@@
Given $(\mathscr{E},\theta)\in{\rm HIG}_{p-1}(\mathscr{X}^{\prime}_{\rm log}/k,G)$, we obtain a logarithmic flat $G$-bundle $(\mathscr{V},\nabla)$ via a logarithmic version of the inverse Cartier transform given in Theorem 4.2. For convenience, we regard $(\mathscr{E},\theta)$ (resp. $(\mathscr{V},\nabla)$) as equivariant logarithmic $G$-Higgs bundles $(E,\theta)$ (resp. logarithmic flat $G$-bundles $(V,\nabla)$) on ${\rm Spec}(k[w_{1},\dots,w_{d}])$. Thus, following the construction of (3.2) in the proof of Theorem 3.1, we have
@@ 4.4. Stability Condition
In this subsection, we make a further assumption on the root stack $\mathscr{X}\cong[Y/\Gamma]$ that it is projective, i.e. the coarse moduli space $X$ is projective. With respect to a given ample line bundle $H$ on $X$, the degree of a locally free sheaf $\mathscr{F}$ on is given in [Bor07, 4.1.2] and denote it by $\deg\mathscr{F}:=\deg_{H}\mathscr{F}$. Moreover, a big open substack $\mathscr{X}^{\rm big}\subseteq\mathscr{X}$ can be regarded as a $\Gamma$-invariant big open subset $Y^{\rm big}\subseteq Y$ such that $\mathscr{X}^{\rm big}\cong[Y^{\rm big}/\Gamma]$. Then we define stability conditions of logarithmic $\lambda$-connections on $\mathscr{X}$ in the same way as Definition 2.5.
@@
A logarithmic $\lambda$-connection $(\mathscr{E},\nabla)$ on $\mathscr{X}$ is $R$-semistable (resp. $R$-stable), if
  • for any maximal parabolic subgroup $P\subseteq G$;
  • for any antidominant character $\chi:P\rightarrow\mathbb{G}_{m}$ trivial on the center;
  • for any big open substack $\mathscr{X}^{\rm big}\subseteq\mathscr{X}$;
  • for any $\nabla$-compatible reduction of structure group $\sigma:\mathscr{X}^{\rm big}\rightarrow(\mathscr{E}|_{\mathscr{X}^{\rm big}})/P$,
@@
We introduce the following categories
  • ${\rm HIG}^{s(s)}_{p-1}(\mathscr{X}_{\rm log}/k,G)$: the category of $R$-(semi)stable nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$;
  • ${\rm MIC}^{(s)s}_{p-1}(\mathscr{X}_{\rm log}/k,G)$: the category of $R$-(semi)stable nilpotent logarithmic $G$-Higgs bundles of exponent $\leq p-1$ on $\mathscr{X}$.